### v1

# Understanding the Nature of Depth-1 Equivariant Quantum Circuit

Abstract
The Equivariant Quantum Circuit (EQC) for the Travelling Salesman Problem (TSP) has been shown to achieve near-optimal performance in solving small TSP problems ($\leq$ 20 nodes) using only two parameters at depth 1. However, extending EQCs to larger TSP problem sizes remains challenging due to the exponential time and memory for quantum circuit simulation, as well as increasing noise and decoherence when running on actual quantum hardware. In this work, we propose the Size-Invariant Grid Search (SIGS), an efficient training optimization for Quantum Reinforcement Learning (QRL), and use it to simulate the outputs of a trained Depth-1 EQC up to 350-node TSP instances—well beyond previously tractable limits. At TSP with 100 nodes, we reduce total simulation times by 96.4%, when comparing to RL simulations with the analytical expression (151 minutes using RL to under 6 minutes using SIGS on TSP-100), while achieving a mean optimality gap within 0.005 of the RL trained model on the test set. SIGS provides a practical benchmarking tool for the QRL community, allowing us to efficiently analyze the performance of QRL algorithms on larger problem sizes. We provide a theoretical explanation for SIGS called the Size-Invariant Properties that goes beyond the concept of equivariance discussed in prior literature.
1 Introduction
Quantum reinforcement learning (QRL) studies parameterized quantum circuits (PQCs) as function approximators within deep reinforcement learning (DRL) pipelines. A common instantiation replaces the neural network in a Deep Q-Network (DQN) with a PQC trained end to end by a classical optimizer, yielding a hybrid agent. Early demonstrations show that PQC-based DQNs can learn near-optimal policies on small benchmarks (e.g., FrozenLake, Cognitive Radio), establishing feasibility and hinting at alternative inductive biases relative to classical value approximators [chen-qrl-2020].
This paper is concerned with solving combinatorial optimization problems with QRL. In the classical domain, Neural Combinatorial Optimization (NCO) develops DRL/attention architectures that construct high-quality solutions to discrete problems, achieving near-optimal tours for symmetric TSP up to 100 nodes [Kool2019Attention]; more recent diffusion/backbone advances further improve the quality–speed trade-off [Li2024FastT2T].
For QRL to match or even surpass NCO counterparts, two gaps remain. Interpretability. PQC architectures with far fewer parameters than classical networks can perform well, yet the role of each parameter is often opaque for concrete combinatorial tasks. Scalability. Prevalent encodings map one decision variable to one qubit and rely on highly entangling layers; both simulation and hardware execution grow costly with problem size, even with efficient tensor-network simulators. These gaps motivate the developments in this paper.
1.1 Overview of the Literature
QRL for Combinatorial Optimization. QRL involves replacing the value function approximator in Deep Reinforcement Learning (DRL) with a Parameterized Quantum Circuit (PQC). Skolik et al. (2023) marked the first systematic application of QRL to Combinatorial Optimization via the Equivariant Quantum Circuit (EQC) for TSP [Skolik2022EquivariantQC]. Using a Symmetry Preserving Ansatz (SPA), they demonstrated that respecting graph isomorphisms can drastically reduce the number of trainable parameters of the circuit down to two values, while producing near-optimal tours for instances of up to 20 nodes. This also serves as a starting point of our work. More recently, Kruse et al. (2024) generalized this framework to Quadratic Unconstrained Binary Optimization (QUBO) formulations, with an sge-sgv ansatz that closely related to the EQC (having also two parameters per layer), demonstrating the ability to learn near-optimal policies on other Combinatorial Optimization problems such as Weighted-MaxCut, Knapsack and the Unit Containment Problem [Kruse_2024_HEA].
Ansatz design and trainability. An ansatz specifies the a certain structure of a PQC to prepare a variational state or wavefunction [Tilly2022]. The structure of the ansatz can be chosen with or without prior knowledge of the problem. In QRL, the most popular ansatz used is the Hardware Efficient Ansatz (HEA) [chen-qrl-2020, Kruse_2024_HEA, dragan2022qrlHEA] due to their device-friendly layouts. However, due to its high expressivity and problem independence, these ansatz can trigger Barren Plateaus (BP), making gradients vanish as problem sizes increase [BP_Holmes_2022, BP_McClean_2018]. This has fueled a shift toward problem-tailoned ansatzes for QRL (to be symmetry-aware) in order to improve trainability at shallow depths.
We can interpret the work of Skolik et al. (2023) in three ways. Firstly, from a Geometric Quantum Machine Learning (GQML) perspective, Schatzki et al. (2024) formalizes SPAs that are equivariant under the symmetry group $S_{n}$, and prove that $S_{n}$-equivariant ansatzes do not suffer from barren plateaus, quickly reach over-parameterization, and generalize well from small amounts of data [Schatzki_2024, Larocca_2023]. Both the EQC and the sge-sgv ansatzes are examples of $S_{n}$-equivariant Quantum Neural Networks (QNNs). Secondly, Skolik et al. (2023) further empirically demonstrates that SPAs outperform their non-symmetry preserving counterparts through ablation simulations which gradually break the equivariance of the EQC [Skolik2022EquivariantQC]. An alternate lens for interpreting the EQC is via the Quantum Approximate Optimization Algorithm (QAOA). The structure of the EQC mirrors the structure of the QAOA ansatz: where the parameters $\gamma$ and $\beta$ play analogous roles to the cost and mixer Hamiltonians [farhi2014quantumapproximateoptimizationalgorithm]. QAOA is well-known in solving combinatorial optimization problems, prominently on the MaxCut problem with proven bounds [Wurtz_2021, Marwaha_2021] and empirical benchmarks [leo2020, Lotshaw_2021]. Lastly, the idea of ensuring that embeddings of graphs to be invariant under node permutations closely follows to the design principles of classical NCO algorithms that respect graph structures through message passing and attention [gillman2024combiningreinforcementlearningtensor, kaempfer2019learningmultipletravelingsalesmen, kool2019attentionlearnsolverouting].
1.2 Contributions
Our work makes the following contributions:
  • A size-invariant understanding of parameters: We provide a new explanation for the roles of the trainable parameters ($\gamma$ and $\beta$) at Depth 1 of the EQC, extending beyond the concept of equivariance in [Skolik2022EquivariantQC]. By analyzing the closed-form expressions of the Depth 1 EQC [akshay2021parameter, analytical-qaoa, exp-qaoa-ising, qaoa-ferm], we show how these parameters govern effective QRL policies for TSP. On one hand, while QAOA uses variational optimization to minimize Hamiltonian expectations, QRL (using Q-Learning) involves optimizing the Temporal-Difference Loss [sutton-barto-2018], with Q-values derived from ZZZZ expectations and guided by a Markov Decision Process(MDP). This fundamental methodological difference motivates our analysis of Depth 1 EQCs in reinforcement learning.
  • Size-invariant trainability:. As a consequence of the size-invariant interpretation of parameters, we conclude that effective QRL policies with Depth 1 EQCs reside in a compact, problem size-invariant parameter region. Leveraging this, we propose the Size-Invariant Grid Search (SIGS), which reduces training time for EQC-based QRL by 27 times (e.g., 151 minutes (using the Analytical Expression) to 5.43 minutes for TSP with 100 nodes), and enables efficient simulation of instances with up to 350 nodes (longest run completes under 140140 minutes). This allows us to analyze scalability of Depth 1 EQCs well beyond the tractable limits of classical quantum simulation.
Limitation. Our findings are restricted to Depth 1 EQCs. While, in theory, deeper circuits should improve the agent’s performance, our empirical results at EQC Depths 2 to 4 only increases the time cost of simulation, with negligible difference in performance when compared to Depth 1 (see Section LABEL:sect:discussion). This negligible difference in performance is consistent with prior Depth-4 results in [Skolik2022EquivariantQC].
The remainder of the manuscript proceeds as follows: Section 2 outlines the existing methodology in Skolik et al. (2023)[Skolik2022EquivariantQC] which this project builds upon. Section LABEL:sect:size-invariant-theory then outlines the Size-Invariant Properties of the EQC at Depth 1, showing that good QRL policies are found within a constrained search space of parameters. Section LABEL:sect:size-invariant-method presents the Size-Invariant Grid-Search (SIGS) procedure. Finally, Section LABEL:sect:size-invariant-exp empirically validates our theorems through various simulation results.
2 Background and Existing Methodology
In this section, we review the relevant background and EQC-based methodology introduced in [Skolik2022EquivariantQC], providing the necessary context for the contributions presented in the subsequent sections.
2.1 Markov Decision Process for TSP
The formulation of TSP as a Markov Decision Process (MDP) used in Skolik et al. (2023) [Skolik2022EquivariantQC] is described here. We follow the same MDP in this work.
Given a TSP size $|V|$, with $|E|=\frac{|V|(|V|-1)}{2}$ edges, each component of the MDP is defined as follows:
  1. 1.
    State Space: Each (Reinforcement Learning) state is defined as the tuple $(s,d,e,t)$. $s\in\{0,\pi\}^{|V|}$ denotes a one-hot encoding for which nodes have been visited in a graph. The element $s_{i}$ is set to $\pi$ for unexplored nodes, and 0 for explored nodes. $d\in\mathbb{R}^{|E|}$ denotes a vector of pairwise Euclidean Distance, and $e$ is a vector of scaled edge weights $e=\tan^{-1}(d)\in\mathbb{R}^{|E|}$, following the reference implementation in [Skolik2022EquivariantQC]. Note that $f(x)=\tan^{-1}(x)$ is a monotonically increasing function in edge distance $x$. $t\in[1,|V|]$ denotes the current node where the agent resides.
  2. 2.
    Action Space: Each action $a$ is a node index (ie. $1\leq a\leq|V|$, $s_{a}=\pi$). We restrict possible values of $a$ to nodes that have not been explored. The environment then appends this node to the end of the tour.
  3. 3.
    Reward: At each step, the reward is the negative increase in the tour length, measured by Euclidean Distance. Suppose the tour length at step $t$ is $L_{t}$ and the tour length at step $t+1$ is $L_{t+1}$. Then the reward to the agent is $L_{t}-L_{t+1}$. We always assume that the last node of the partial tour connects to the first node, and factor in this edge weight accordingly.
  4. 4.
    Transition Probabilities: In a natural TSP formulation as an MDP, taking an action logically results in a deterministic next state: the node $a$ becomes explored, and $s_{a}\leftarrow 0$ with probability 1.
  5. 5.
    Episodes and Steps: An episode represents a complete TSP tour. The initial step $t=0$ is the initial state the RL agent is in, denoted by $s_{0}$. Without loss of generality, our tours begin with the node at index 11 (assuming one-indexed), where only node 1 is visited. The terminal step $T$ is the step |V|1|V|-1, which is when all nodes are visited, and $t$ being some permutation of the $|V|$ vertices.
2.2 Parameterized Quantum Circuit
The Equivariant Quantum Circuit (EQC) is an example of an PQC. A diagram of the EQC is provided in Figure LABEL:fig:eqc-diagram for a $n=4$ qubit system. In the formulation of the EQC, there is a one-to-one mapping between qubits and TSP nodes, which means an $n$ qubit system is only able to handle TSP instances of size $n$. There are two kinds of parameterized gates in the EQC as shown: a ZZZZ-gate and a $R_{x}$ gate. Note that $\gamma$ is shared across all ZZZZ-gates, while $\beta$ is shared across all $R_{x}$ gates. Measurements are recorded with the $ZZ=Z\bigotimes Z$ observable. The analytical expression for the output of the EQC in the basis of the observable $\langle Z_{t}Z_{a}\rangle$, where $t$ is the last node of the present tour and $a$ is an unexplored next node, is given in [Skolik2022EquivariantQC]:
$\langle Z_{t}Z_{a};(\gamma,\beta)\rangle=\sin(\pi\beta)\sin(\gamma e_{ta})\prod_{k=1,k\neq t}^{n}\cos(\gamma e_{ak}).$ (1)